A wheel of mass 6.0kg and radius of gyration 40cm is rotating at 300prm. Find it's moment of inertia and it's rotational KE
1
Expert's answer
2020-05-25T11:20:06-0400

By definition, the moment of inertia of the body with the radius of gyration "R_g = 40 cm" and mass "m = 6kg" will be:


"I = mR_g^2 = 6\cdot 0.4^2 = 0.96 kg\cdot m^2"


The rotational KE:


"E_{rot} = \dfrac12I\omega^2"

where "\omega" is the angular velocity. If the frequency of rotation is "\nu = 300 rpm = 5Hz", then the angular velocity will be:


"\omega = 2\pi\nu = 2\pi\cdot 5 \approx 31.4 rad/s"

Thus, the rotational KE will be:


"E_{rot} = \dfrac12\cdot 0.96\cdot 31.4^2 = 473.7 J"

Answer. I = 0.96 kg*m^2, Erot = 473.7 J.